OpenStudy (anonymous):

solve the rational equation \[\[\frac{ m }{ 6 }-\frac{ 5 }{ 3 }=\frac{ 15 }{ m^2-25}\] \]

OpenStudy (anonymous):

under the 15 should be m+5 m_5

OpenStudy (jdoe0001):

$$\bf \frac{ m }{ 6 }-\frac{ 5 }{ 3 }=\frac{ 15 }{ m^2-25}\\ \frac{m-10}{6} = \frac{ 15 }{ m^2-25} $$ then just cross multiply

OpenStudy (jdoe0001):

I get \(\bf m^3-10m^2-25m-250=80\\ m^3-10m^2-25m-330 = 0\)

OpenStudy (anonymous):

My first comment was wrong

OpenStudy (jdoe0001):

\(\bf \cfrac{ m }{ 6 }-\cfrac{ 5 }{ 3 }=\cfrac{ 15 }{ (m+5)(m-5)}\ \ ?\)

OpenStudy (anonymous):

no I mixed two problems up on accident \[\frac{ m }{ 6 }-\frac{ 5 }{ 3 }=\frac{ x+2 }{ 3}\]

OpenStudy (anonymous):

Thats the real problem^^

OpenStudy (jdoe0001):

same thing \(\bf \large \cfrac{ m }{ 6 }-\cfrac{ 5 }{ 3 }=\cfrac{ x+2 }{ 3}\\ \cfrac{m-10}{6} = \cfrac{ x+2 }{ 3}\) then cross multiply

OpenStudy (jdoe0001):

6x-3m +32 = 0 is what I get

OpenStudy (jdoe0001):

erk

OpenStudy (jdoe0001):

6x-3m +42 = 0 is what I get

OpenStudy (anonymous):

oh okay thanks :)

OpenStudy (jdoe0001):

yw